OpenStudy (loser66):

Find the solution of the given initial value problem in explicit form y'=(1-2x)/y ; y(1)=-2

OpenStudy (loser66):

\[\frac{dy}{dx}y =1-2x\\ydy=(1-2x)dx\] integral both sides I got the answer is y =+-sqrt(2x-2x^2 +c solve for c , I have c=4 so the answer should be +- sqrt(2x+2x^2+4 However, the answer in book reject positive answer, why?

OpenStudy (phi):

are you saying the answer is \[ y= \pm \sqrt{2x - 2x^2 +4} \] at x=0 you would get \[ y = \pm\sqrt{4} = \pm 2\] based on y(1)= -2, you can only use the - sign solution

OpenStudy (dan815):